Solve the system by addition or substitution. 4x – y = –3 y = 4x + 8
Answer by jai_kos(139)

Solution:
Given 4x – y = –3 y = 4x + 8
Let us divide the equation in two, we get
4x – y = -3y --------(1)
And
-3y = 4x + 8 --------(2)
Consider the equation(1), we get
4x – y = - 3y
Put the entire y in one side, we get
4x = -3y + y
4x = -2y
-2x = y --------(3)
Now substitute equation(3) in equation(2), we get
-3(-2x) = 4x + 8
6x = 4x + 8
Get the entire “x” on one side, we get
6x - 4x = 8
2x = 8
x = (8/2)
x = 4 ---------(4)
Now put the value of “x” in equation(3), weget
-2(4) = y
-8 = y
Hence the value of “x” = 4 and the value of “y” = -8